# InstrAug: Automatic Instruction Augmentation for Multimodal Instruction Fine-tuning

### v1

@@ 4.4 Scaling Factors
We also notice the scaling factors from the perspective of the dataset (including instruction and instance) and model size.
@@ 5.1 Potential Impact Factors to MIns+
As mentioned above, InstrAug can improve the performance on the benchmark level. Still, we are curious about what dominates and how much each individual task can benefit from instruction augmentation. To probe the potential factors at the task level, we define the following task particulars:
@@
We display the average gain for each task on OFA (59K) together with the individual attributes in Figure 3. The Pearson Correlation Coefficient between instance text proportion and the average performance gain is 0.639, which uncovers a strong correlation between the relative length of instance text to the extra gain by InstrAug. Moreover, tasks with shorter instance text fall behind those with longer instance text. A probably reasonable explanation for this is that MLLMs attend excessively to the task-irrelevant templates, which could perturb the gradient to an inappropriate direction and scale Von Oswald et al. (2023).
When the problem to solve has been entirely enclosed in the template like VTE, the task can benefit more from InstrAug, since for such tasks in the training set InstrAug is imposed directly on the question rather than modifier text. Compared to those whose instructions contain options after instantiation, tasks without visible options (TextVQA and VisDial) are susceptible to deterioration on MIns+, possibly because instruction augmentation can help consolidate the copy-from-input ability by modifying the surrounding context, while pure generative tasks do not rely on this ability to predict answers.
@@ 6 Conclusion
In this paper, we propose InstrAug, a completely automatic instruction augmentation framework for large instruction-following datasets with very scarce hand-written instructions. The framework purely relies on an arbitrary LLM and simple meta-prompts. We specially devise the generation process and rule-based filters based on the characteristics of raw instructions and model outputs. Experimental results on two MLLMs across a broad range of tasks demonstrate our proposed method can significantly enhance the overall performance compared to baselines using non-augmented datasets, which also shows the improved robustness of human-written instructions in different wording styles.
@@
In InstrAug, we apply a series of rule-based filters to remove invalid outputs that could do evil to the dataset construction and tuning process, as well as a heuristic sampling strategy to balance the distribution of generated instructions in different versions of MIns+. However, InstrAug works on augmentation with less focus on other dataset surgery techniques such as data selection Xia et al. (2024) and dataset pruning Yang et al. (2022), which we believe may further improve the current performance by combining them with InstrAug. We hold this thought as our future research direction.
@@
We clarify ethical considerations in two aspects. First, InstrAug is a fully automatic augmentation framework and no crowdworker is hired. Second, regarding the generated process from LLM (LLaMA-2-13b-Chat), since it is an open-sourced model with safety-tuning, and our instructions do not contain question or instructions that deliberately elicit harmful output from the LLM. We have thoroughly checked the augmented instruction set—apart from possible hallucinations, most of the output follows the guiding instruction and just modifies the extant instructions in MIns. Moreover, we provide reference inputs and outputs in the appendix to make the generation process transparent and public, and commit to release the instruction set for public inspection.
Appendix A More Experimental Results
@@
Appendix B Generation Details
B.1 Guiding Instruction Generation
@@ B.2 New Instruction Generation
B.3 The Intuition behind Length Filter
@@
[System Prompt] |
You are a helpful assistant. |
[User] |
Generate 10 instructions about how to rephrase short text. Your response should be cononical and formatted as enumerations. |
[System] |
@@
Use synonyms (1) | Sentiment Understanding | Raw |
What is the caption of the image? |
Gen |
What is the description of the image? |
Object Region Match | Raw |
Is the object {text} in {regions}? {options} |
Gen |
Is the object {text} located in {regions}? {options} |
Rearranging Phrases (2) | Image Caption | Raw |
Please analyze the sentiment depicted in the image and answer the question. {question} {options} |
Gen |
Please answer the question after analyzing the sentiment depicted in the image. {question} {options} |
Object Region Selection | Raw |
In this task, you are given the coordinates of some rectangular region in the image. You need to first localize each rectangular region and then identify what is the object in the region. The region is {regions}. |
Gen |
The task at hand is to localize and identify objects within a specified rectangular region in an image. The region of interest is {regions}. To complete the task, you will need to first localize the region and then identify the object within it. |
Simplyfy Language (3) | Grounded Captioning | Raw |
What is the content of {regions}? |
Gen |
What is inside {regions}? |
Visual Attribute | Raw |
Decide which option is the attribute of the object in the given region. Region: {regions} {options} |
Gen |
Determine the attribute of the object in the specified region: {regions} {options} |
Remove Redundancy (4) | Image Caption | Raw |
In this task, you will look at the image and briefly describe the image. |
Gen |
Briefly describe the image you see. |
Object Grounding | Raw |
Given the description of an object, generate the region that contains this object. The description is: {text} |
Gen |
Please generate the region that contains the object described by ’{text}’. |
Use Active Voice (5) | Factual Checking | Raw |
Deicide if the claim can be supported by the image and the context. Context: {context} Claim: "{text}" {options} |
Gen |
Evaluate the validity of the claim {text} {options} based on the image and the surrounding context {context} . |
Object Identification | Raw |
In this task, you are required to answer a question about the appearance of an object.{question} {options}. |
Gen |
For this task, you need to describe the appearance of an object. Please provide your answer in the format of {question} {options}. |
Appendix C MIns+ Details
C.1 Statistics of MIns+
@@
C.2 Examples of Rewritten Instructions
@@ C.4 Wording Characteristics
Appendix D Training and Evaluation Details
D.1 Hardware and Hyperparameters
@@ D.2 Tuning Details
D.3 More Details of Evaluation Tasks
@@
VizWiz | MC |
D.4 Evaluation Methods
@@
D.5 Templates for IBLIP-Bench Tasks
We provide the evaluation templates for two multiple choice IBLIP-Bench tasks (Science QA, Icon QA) in Table 11. The raw instruction can be found in Dai et al. (2023), which we show in the row “IBLIP”. The instructions in the “OFA” row are what we create from the raw instruction by modifying delimiters and special tokens to be compatible with OFA’s expression style. Here to simplify the expression, we assume there are two options, “Apple” and “Pear” and the correct answer is always “Apple”. We also list the target strings to judge whether the output is correct.
@@ D.5 Templates for IBLIP-Bench Tasks
Appendix E Result Visualization
@@
![[Uncaptioned image]](x7.png)
![[Uncaptioned image]](x8.jpg)

### v2

@@ 4.3 Scaling Laws among Results
We also notice the scaling laws as mentioned in Hoffmann et al. (2022) in terms of the dataset (including instruction set and number of samples) and model size.
@@ 5.1 Potential Impact Factors to MIns+
As mentioned above, InstrExp can improve the performance on the benchmark level. Still, we are curious about what dominates and how much each individual task can benefit from instruction augmentation. To probe the potential factors at the task level, we define the following task particulars:
@@
We display the average gain for each task on OFA (59K) together with the individual attributes in Figure 3. The Pearson Correlation Coefficient between instance text proportion and the average performance gain is 0.639, which uncovers a strong correlation between the relative length of instance text to the extra gain by InstrExp. Moreover, tasks with shorter instance text fall behind those with longer instance text. A probably reasonable explanation for this is that MLLMs attend excessively to the task-irrelevant templates, which could perturb the gradient to an inappropriate direction and scale Von Oswald et al. (2023).
When the problem to solve has been entirely enclosed in the template like VTE, the task can benefit more from InstrExp, since for such tasks in the training set InstrExp is imposed directly on the question rather than modifier text. Compared to those whose instructions contain options after instantiation, tasks without visible options (TextVQA and VisDial) are susceptible to deterioration on MIns+, possibly because instruction augmentation can help consolidate the copy-from-input ability by modifying the surrounding context, while pure generative tasks do not rely on this ability to predict answers.
@@ 6 Conclusion
In this paper, we propose InstrExp, a completely automatic instruction augmentation framework for large instruction-following datasets with very scarce hand-written instructions. The framework purely relies on an arbitrary LLM and simple meta-prompts. We specially devise the generation process and rule-based filters based on the characteristics of raw instructions and model outputs. Experimental results on two MLLMs across a broad range of tasks demonstrate our proposed method can significantly enhance the overall performance compared to baselines using non-augmented datasets, which also shows the improved robustness of human-written instructions in different wording styles.
@@
In InstrExp, we apply a series of rule-based filters to remove invalid outputs that could do evil to the dataset construction and tuning process, as well as a heuristic sampling strategy to balance the distribution of generated instructions in different versions of MIns+. However, InstrExp works on augmentation with less focus on other dataset surgery techniques such as data selection Xia et al. (2024) and dataset pruning Yang et al. (2022), which we believe may further improve the current performance by combining them with InstrExp. We hold this thought as our future research direction.
@@
We clarify ethical considerations in two aspects. First, InstrExp is a fully automatic augmentation framework and no crowdworker is hired. Second, regarding the generated process from LLM (LLaMA-2-13b-Chat and GPT-4), since they have experienced sufficient safety-tuning, and our instructions do not include deliberate attack to elicit harmful output, we believe there is extremely low risk in generating harmful information. We have thoroughly checked the augmented instruction set—apart from possible hallucinations, most of the output follows the guiding instruction and just modifies the extant instructions in MIns. Moreover, we provide reference inputs and outputs in the appendix to make the generation process transparent and public, and commit to release the instruction set for public inspection.
Appendix A Baseline Models
In this section, we elaborate on the details of the models evaluated in Section 4.
OFA
InstructBLIP
LLaVA
Liu et al. (2023b, a) LLaVA represents a novel end-to-end trained large multimodal model. It connects vision encoder to a pretrained LLM Vicuna for general-purpose visual and language understanding. There are total two training phases to produce LLaVA: 1) Feature alignment (pretraining): this stage aligns the visual features output from the vision encoder to the language’s hidden space. The pre-training is based on 558K subset of LAION-CC-SBU dataset Schuhmann et al. (2021) with both LLM and visual encoder frozen. 2) End-to-end tuning: GPT-4 is employed to provide synthetic multimodal dialogue data, which is then tuned along with science QA benchmark to further enhance the instruction-following ability of LLaVA. It achieves impressive chat capabilities reminiscent of the multimodal GPT-4 and sets a new state-of-the-art accuracy on Science QA.
Due to limited computation resources, we use LoRA Hu et al. (2021) to finetune LLaVA-vicuna-7B-v1.6 on our generated instruction dataset.
Appendix B More Experimental Results
@@
Appendix C Generation Details
C.1 Guiding Instruction Generation
@@ C.2 New Instruction Generation
C.3 The Intuition behind Length Filter
@@
| [System Prompt] | You are a helpful assistant. |
| [User] | Generate 10 instructions about how to rephrase short text. Your response should be cononical and formatted as enumerations. |
| [System] | (Irrelevant output) 1. Use synonyms: Replace specific words or phrases with their synonyms to convey the same meaning for the input. 2. Rearrange phrases for the input sentence: Reorganize sentences or phrases to improve readability or flow without changing the overall meaning. 3. Simplify language: Use simpler vocabulary and sentence structures to make the text more accessible and easier to understand. 4. Remove redundancies: Eliminate unnecessary words or phrases that do not add to the meaning of the text. 5. Use active voice: Convert passive sentences to active ones to make the text more engaging and easier to read. 6. Break up long sentences: Divide long sentences into shorter ones to improve readability and flow. 7. Use bullet points or numbered lists: Organize information using bullet points or numbered lists to make it easier to read and understand. 8. Replace jargon with common terms: Use plain language instead of technical jargon or industry-specific terms to make the text more accessible to a wider audience. 9. Use inclusive language: Avoid using language that stereotypes or excludes certain groups of people. Instead, use language that is respectful and inclusive of all individuals. 10. Check for consistency: Ensure that the text is consistent in terms of tone, style, and format throughout. (Irrelevant output) |
@@
| Use synonyms (1) | Sentiment Understanding | Raw | What is the caption of the image? |
| Gen | What is the description of the image? | ||
| Object Region Match | Raw | Is the object {text} in {regions}? {options} | |
| Gen | Is the object {text} located in {regions}? {options} | ||
| Rearranging Phrases (2) | Image Caption | Raw | Please analyze the sentiment depicted in the image and answer the question. {question} {options} |
| Gen | Please answer the question after analyzing the sentiment depicted in the image. {question} {options} | ||
| Object Region Selection | Raw | In this task, you are given the coordinates of some rectangular region in the image. You need to first localize each rectangular region and then identify what is the object in the region. The region is {regions}. | |
| Gen | The task at hand is to localize and identify objects within a specified rectangular region in an image. The region of interest is {regions}. To complete the task, you will need to first localize the region and then identify the object within it. | ||
| Simplyfy Language (3) | Grounded Captioning | Raw | What is the content of {regions}? |
| Gen | What is inside {regions}? | ||
| Visual Attribute | Raw | Decide which option is the attribute of the object in the given region. Region: {regions} {options} | |
| Gen | Determine the attribute of the object in the specified region: {regions} {options} | ||
| Remove Redundancy (4) | Image Caption | Raw | In this task, you will look at the image and briefly describe the image. |
| Gen | Briefly describe the image you see. | ||
| Object Grounding | Raw | Given the description of an object, generate the region that contains this object. The description is: {text} | |
| Gen | Please generate the region that contains the object described by ’{text}’. | ||
| Use Active Voice (5) | Factual Checking | Raw | Deicide if the claim can be supported by the image and the context. Context: {context} Claim: "{text}" {options} |
| Gen | Evaluate the validity of the claim {text} {options} based on the image and the surrounding context {context} . | ||
| Object Identification | Raw | In this task, you are required to answer a question about the appearance of an object.{question} {options}. | |
| Gen | For this task, you need to describe the appearance of an object. Please provide your answer in the format of {question} {options}. |
Appendix D MIns+ Details
D.1 Statistics of MIns+
@@
D.2 Examples of Rewritten Instructions
@@ D.4 Wording Characteristics
Appendix E Training and Evaluation Details
E.1 Hardware and Hyperparameters
@@ E.2 Tuning Details
E.3 More Details of Evaluation Tasks
@@
| VizWiz | MC | |||||
E.4 Evaluation Methods
@@
E.5 Templates for IBLIP-Bench Tasks
We provide the evaluation templates for three IBLIP-Bench tasks (Science QA, Icon QA, VizWiz) in Table 13. The raw instruction can be found in Dai et al. (2023), which we show in the row “IBLIP”. The instructions in the “OFA” row are what we create from the raw instruction by modifying delimiters and special tokens to be compatible with OFA’s expression style. Here to simplify the expression, we assume there are two options, “Apple” and “Pear” and the correct answer is always “Apple”. We also list the target strings to judge whether the output is correct.
@@ E.5 Templates for IBLIP-Bench Tasks
Appendix F Result Visualization
@@
| Evaluation result example 1 (Generation): Visual Text Extraction | |
| Input Image |
| Target | look at her… she didn’t learn anything from the last time |
| Instruction 1 | What is the text written on the image? |
| MIns (59K) | a woman holding a surfboard |
| MIns+ (59K) | look at her, she didn’t learn anything from the last time |
| Instruction 2 | This image contains some text. For this task, you need to look at the image carefully and identify all the text in the image. The text in the image is |
| MIns (564K) | yes |
| (564K) | she didn’t learn anything from the last time |
| Evaluation result example 2 (Multiple Choice): Disaster Classification | |
| Input Image |
| Target | not disaster |
| Instruction 1 | What disaster happens in the image? [Options]: earthquake||||informative||||not disaster||||mild damage||||other disaster||||hurricane||||little or none damage||||not informative||||severe damage||||landslide||||flood||||fire |
| MIns (59K) | earthquake |
| MIns+ (59K) | not disaster |
| Instruction 2 | According to the image, what kind of disaster happened? Choose the correct answer from options. [Options]: earthquake||||informative||||not disaster||||mild damage||||other disaster||||hurricane||||little or none damage||||not informative||||severe damage||||landslide||||flood||||fire |
| MIns (564K) | hurricane |
| (564K) | not disaster |
![[Uncaptioned image]](extracted/5606230/figure/vte_case.png)
![[Uncaptioned image]](extracted/5606230/figure/dc_case.jpg)